# Protocol Futuring: Speculating Second-Order Dynamics of Protocols in Sociotechnical Infrastructural Futures

### v1

@@ 2.1.4. Speculative and Fictional Protocols in Design and HCI
These experiments treat protocols not only as technical artifacts but also as sites for critical imagination. By dramatizing potential breakdowns or alternate futures, they surface hidden assumptions, ethical dilemmas, and opportunities for resilience that conventional planning might miss.
@@ 2.2. Design Futuring
Speculative design and design fiction have emerged over the past two decades as important methods in HCI and design research for interrogating future possibilities. These approaches create fictional artifacts, scenarios, or narratives that encourage audiences to question assumptions and consider alternative social or technological trajectories. As originally articulated by designers like Dunne & Raby, speculative design emphasizes provoking critical reflection on social values and ethics by presenting uncanny or exaggerated design concepts (often as physical prototypes or visual media) (dunne2013speculative). Design fiction, a term popularized by science fiction author Bruce Sterling and designer Julian Bleecker, is closely related: it involves using fictional diegetic prototypes—artifacts that exist within an imagined future world—to “suspend disbelief about change” and spark discussion (sterling2005shaping; bleecker2009design). Sterling notably described design fiction as “the deliberate use of diegetic prototypes to suspend disbelief about change”, suggesting that its power lies in focusing people on a specific object or service in context, rather than on abstract futurism or grand policy. In practice, design fictions often take the form of videos, stories, or concept designs that ask “what if…?” about near-future technologies—for example, envisioning new gadgets and their everyday use, or hypothetical tech policies and their societal impact.
@@
Participatory design (PD), on the other hand, originates from a different tradition (Scandinavian cooperative design in the 1970s) and focuses on involving stakeholders—especially end users or marginalized groups—directly in the design process. The core idea of PD is to democratize design decision-making and ensure outcomes reflect the needs and values of those affected (simonsen2012routledge). While classic PD was applied to developing real systems, there has been growing interest in participatory speculative design: inviting participants to co-create visions of the future. This fusion aims to overcome one criticism of early speculative design and design fiction: that they sometimes reflected the narrow, “privileged” perspectives of their authors (tonkinwise2014design). Scholars have pointed out that many speculative design works, especially from Western academic or art contexts, carried assumptions not shared by broader populations, and risked a kind of cultural or technological elitism. In response, methods like participatory design fiction (e.g. workshops where diverse groups build fictional scenarios together) and speculative co-design have been developed to inject plurality and ground speculation in a wider range of experiences (lindley2017anticipatory; gerber2020participatory). For instance, researchers have used speculative methods in community settings to envision futures of aging, healthcare, or urban living, treating participants as “world-builders” of their own futures rather than just consumers of a designer’s vision (Kozubaev2020Expandingb).
@@
Another notable evolution in design fiction practice has been a shift from focusing on single artifacts to considering systems and infrastructures. Sterling’s notion of keeping design fiction “on the level of objects and services” (rather than entire geopolitical shifts) was a useful starting heuristic, but many contemporary design researchers have sought to reconnect those isolated artifacts to the larger contexts that give them meaning. As Wong et al. observe, speculative designs “often imply a lifeworld” even if they only show a gadget (wong2020infrastructural). There is an increasing push to explicate that lifeworld: to design not just a speculative object but also illustrate the networks, practices, regulations, and cultures around it. Infrastructural Speculations, as introduced by Wong, exemplify this trend. They outline design tactics to deliberately incorporate infrastructure considerations into speculative design, so that the speculation “places lifeworlds at the center of design concern” and highlights the cultural, regulatory, environmental, and repair conditions enabling the imagined future. Their work suggests techniques like “infrastructural inversion”—essentially flipping the focus to the normally invisible support systems—and considering multiple stakeholders and temporal scales in speculative scenarios.
@@ 2.3. Infrastructure Studies, Temporality, and Sociotechnical Imaginaries
Our work is additionally informed by the rich body of infrastructure studies in both HCI and Science & Technology Studies (STS). Infrastructure in this context refers to the large-scale sociotechnical systems and underlying structures that support human activities—from transportation grids and power networks to information systems and organizational routines. A foundational perspective by Star and Ruhleder reconceptualized infrastructure not as a static thing, but as something that is relational, ecological, and learned through use (Star1996). Bowker et al. argue for “information infrastructure studies” as a way of attending to the mundane yet consequential work of standardization, classification, and data curation that underpins contemporary knowledge practices (bowker2010toward). Edwards and colleagues outline an agenda for infrastructure studies that emphasizes heterogeneity, embeddedness, and the long temporal horizons of infrastructural systems (edwards2003infrastructure). Ribes and Finholt describe the “long now” of infrastructure development, highlighting tensions between innovation and stability, local flexibility and global interoperability, short-term projects and long-term maintenance (Ribes2007Tensions).
A recent systematic literature review by Lyu et al. identifies three major themes in SIGCHI infrastructure research (lyu2025infrastructure):
@@
Subsequent work has elaborated concepts such as infrastructure time and infrastructuring. Karasti, Baker, and Millerand examine how collaborative data infrastructures in science require ongoing care, negotiation, and reconfiguration over years and decades, rather than one-off design interventions (karasti2004infrastructuring). Infrastructuring, drawn from participatory design, foregrounds users and communities as co-designers who continuously shape, appropriate, and re-align infrastructures in practice (Karasti2004; bjorgvinsson2012agonistic). Larkin analyzes “the politics and poetics of infrastructure,” showing how infrastructures materialize state power, cultural imaginaries, and sensory experiences (Larkin2013Politics). Anthropological accounts of infrastructure in cities, energy systems, and logistics further emphasize uneven access, differential vulnerabilities, and the improvisational practices people develop to cope with breakdowns (anand2018promise; Harvey2012Topological).
Crucially for our purposes, STS has also foregrounded temporality and sociotechnical imaginaries as central to infrastructural politics. Jasanoff and Kim conceptualize sociotechnical imaginaries as collectively held, institutionally stabilized visions of desirable futures that guide large-scale techno-political projects (jasanoff2015dreamscapes; Jackson2015). Sareen and Grandin argue that studying energy infrastructures requires explication of temporality, including how expectations, delays, and irreversibilities shape what futures become possible (sareen2021matter). Infrastructure studies thus provide both conceptual tools (e.g., inversion, infrastructuring) and temporal sensibilities (e.g., long-now, path-dependence, legacy) that are highly relevant to protocol design.
@@ 3. Methodological Framework: Protocol Futuring
Protocol futuring is a methodological framework for speculating on the future of sociotechnical infrastructures by using design fiction techniques centered on protocols. In essence, it asks researchers, designers, and participants to imagine future worlds where the key drivers or points of inquiry are protocols—whether technical standards, social rules, or political processes—and to explore the second-order consequences of those protocols in action. This approach extends speculative design into what we call an “infrastructural lifeworld” mode, where the unit of analysis isn’t a single artifact or user experience but rather the network of relationships, agreements, and hidden work that a protocol coordinates or disrupts. This situates our method within, but distinct from, established speculative design traditions (Star1996; bowker1999sorting; pipek2009infrastructuring).
@@
In contrast to previous work, our primary aim is to shift speculative design’s focus from products to protocols. We concentrate on new protocols and how they function in future lifeworlds. Rather than emphasizing specific individuals or heroes in our speculative fiction, we focus on the systemic effects of protocols. Our approach is guided by what we call Chiang’s Law, named after science fiction author Ted Chiang (Rao2025Strange): science fiction is about strange rules, while fantasy is about special people. This aligns the method with recent moves in HCI to examine governance-by-infrastructure, standards, and rule systems (lampland2009standards; ribes2017notes), yet uses fiction to expose how those rules behave over extended temporal horizons.
@@
A primary aim is to move beyond first-order effects (new technology X performs function Y”) to examine second-order and longer-term effects (as X becomes widespread, it causes Z unintended outcome among society/environment, or actors respond with W adaptation). Second-order effects are explored not as deterministic predictions but as plausible dynamics grounded in what Edwards calls the “vast machine” of infrastructural interdependence (edwards2003infrastructure). The method deliberately directs attention to what happens after a protocol’s initial deployment and uptake—essentially, to consider the trajectory of a protocol through time (e.g., years, decades) and scale (from small pilot to global infrastructure). This exploration includes phenomena like protocol drift (implementations or uses diverging from original design over time), protocol jams (congestions or deadlocks in socio-technical systems caused by interacting protocols, similar to traffic jams), and protocol ossification (protocols becoming rigid and resistant to change as they entrench). For example, a Protocol Futuring about a future internet standard might explore how that standard initially improves security but later leads to a new form of digital lock-in or black market to circumvent its restrictions—outcomes beyond the protocol’s immediate goal. By targeting these ripple effects, Protocol Futuring aligns with the ethos of ”predicting the traffic jam, not just the car” (pohl1973), using imaginative narrative as a foresight tool.
@@
In Protocol Futuring, context is not backdrop but a primary analytic unit: scenarios are constructed to show how a protocol operates within—and is reshaped by—the social norms, institutions, regulatory conditions, and legacy infrastructures that constitute its lifeworld. This aligns the method with infrastructural speculation (Wong et al. 2020) and with feminist and care-centered infrastructuring traditions that foreground ongoing maintenance and situated negotiation (pipek2009infrastructuring; karasti2014infrastructuring). Rather than presenting protocols as neutral technical rules, Protocol Futuring depicts how they are interpreted, contested, and maintained over time. This may involve narrative techniques such as following invisible actors (e.g., maintainers, governance committees) or describing scenes from the protocol’s point of view. The aim is to make the infrastructural “wires and threads” visible and to situate protocols within the lived practices that give them force.
@@
Because protocols distribute authority across designers, implementers, users, regulators, adversaries, and bystanders, Protocol Futuring relies on participatory methods to surface these divergent viewpoints. Building on traditions in participatory design (ehn2008participation; bjogvinsson2012design) and game-based co-design (gaver1999design; brandt2012tools), our approach invites multiple contributors to co-construct speculative scenarios through workshops, role-play sessions, and open fiction calls. This distributes authorship beyond designers and enables participants to inhabit different stakeholder positions—everyday users, maintainers, marginalized groups, or adversaries—thereby generating plural, sometimes conflicting, worldviews. Such plurality aligns with the need to imagine protocols that require collective adoption and negotiation rather than heroic individual use. In practice, we draw on PD facilitation techniques (e.g., scenario mapping) and design-game mechanics (brandt2012tools) to structure interaction. These participatory and game-like structures make the infrastructural politics of protocols experientially accessible and foreground how rules shape—and are shaped by—collective imagination.
@@
Although protocols can seem abstract, Protocol Futuring materializes speculative worlds through concrete artifacts, drawing from traditions in critical design (dunne2013speculative), design fiction (sterling2005shaping), and future-oriented prototyping (halse2010rehearsing). Outputs may include short stories, fictional policy documents, speculative diagrams, interface mockups, or immersive videos—forms that make the protocol’s logic, breakdowns, and sociopolitical effects legible. These artifacts function not as “hero objects” but as probes that reveal how a rule system organizes everyday life. For instance, a fabricated citizen handbook for a future data protocol or a visualization of rerouted network flows during a protocol split can foreground operational constraints, power dynamics, and maintenance work. We often employ cross-media storytelling—text, images, interactive maps, AI-generated video—to surface different infrastructural layers. As demonstrated in the South Beast Asia case, such multimodal artifacts help participants inhabit the protocol’s world and expose second-order dynamics that would remain invisible in purely textual speculation. Protocol Futuring thus describes not a fixed output format but a strategic orientation capable of using any medium necessary to render protocol-driven futures thinkable.
@@
Methodologically, we treat Protocol Futuring as a form of research-through-design (RtD) (Blythe2014Research). This means the act of creating the fiction is itself a mode of inquiry: by designing these speculative scenarios, we are effectively probing research questions about protocols (“what if this protocol were in place, how would X or Y play out?”) and generating knowledge in the form of narrative insights, identified pitfalls, or design considerations, following Zimmerman et al (zimmerman2007research). Throughout our process, we maintain a reflexive stance—we document the decision-making, the alternatives considered, the influences from real-world references, and the participants’ reactions. Analysis of a protocol fiction may involve looking for emergent themes across different stories, or using the fiction as a provocation in interviews with experts. For example, after producing a set of stories in the Knowledge Futurama workshop, we conducted a debrief discussion where librarians and archivists (some of whom participated) reflected on the scenarios: which seemed plausible, which touched on real concerns, which surprised them by revealing a blind spot in current practices. This reflective step, aligning with (ghajargar2021synthesis), ensures that the fictional outputs are brought back to bear on present understanding, closing the loop of speculation to inform reality.

### v2

@@ 2.1.4. Speculative and Fictional Protocols
These experiments treat protocols not only as technical artifacts but also as sites for critical imagination. By dramatizing potential breakdowns, alternative coordination mechanisms, or speculative futures, they surface hidden assumptions, ethical dilemmas, and opportunities for resilience that conventional planning might miss—providing a foundation for our own methodological contribution.
@@ 2.2. Design Futuring
Speculative design and design fiction have emerged over the past two decades as important methods in HCI and design research for interrogating future possibilities (Celik and Kaya, 2025). These approaches create fictional artifacts, scenarios, or narratives that encourage audiences to question assumptions and consider alternative social or technological trajectories. As originally articulated by designers like  Dunne and Raby (2013), speculative design emphasizes provoking critical reflection on social values and ethics by presenting uncanny or exaggerated design concepts (often as physical prototypes or visual media). Design fiction, a term popularized by science fiction author Bruce Sterling and designer Julian Bleecker, is closely related: it involves using fictional diegetic prototypes—artifacts that exist within an imagined future world—to “suspend disbelief about change” and spark discussion (Sterling, 2005; Bleecker, 2009). Sterling notably described design fiction as “the deliberate use of diegetic prototypes to suspend disbelief about change”, suggesting that its power lies in focusing people on a specific object or service in context, rather than on abstract futurism or grand policy. In practice, design fictions often take the form of videos, stories, or concept designs that ask “what if…?” about near-future technologies—for example, envisioning new gadgets and their everyday use, or hypothetical tech policies and their societal impact. A recent systematic literature review identifies four core phases common to speculative design processes: select, explore, transform, and provoke (Cardenas Cordova et al., 2025).
2.2.1. Reflective Modes and Critical Perspectives
Kozubaev et al. (2020) contribute five reflective modes to help HCI researchers improve design futuring work: designerly form-giving (the specificity and experiential qualities of design objects), temporality (how futures are represented across time), positionality (recognizing that futuring emerges from particular situated perspectives), real-world engagement (clarifying how speculation connects to actual contexts), and knowledge production (understanding what kinds of knowledge design futuring generates). These modes function as analytical resources for practitioners to articulate, evaluate, and generate new design futuring work, addressing methodological challenges as design futuring becomes more established in HCI.
Howell et al. (2021) critique the Futures Cone—a prominent model in design futuring—for reducing diverse perspectives to a singular point of “the present” and embedding assumptions about linear progress. This work responds to calls for decolonizing futures approaches by acknowledging that there is no single everyday experience and that assuming universality obscures the complexity of historical context and diverse, situated presents.
A recent literature review of HCI’s future-orientation (Sanchez et al., 2025) argues that while HCI is inherently future-oriented, the visions of futures in HCI publications tend to be implicit, techno-deterministic, narrow, and lacking in roadmaps and attention to uncertainties. However, a growing body of HCI research is challenging techno-centric visions by demonstrating active exploration of uncertainty, focus on human experience, and contestation of dominant narratives. Bendor and Lupetti (2024) address speculative design as pedagogical practice, examining how to integrate this “critical material engagement with possible futures” into design-engineering curricula and identifying obstacles for curriculum integration.
2.2.2. Participatory Speculative Design
@@
Participatory design (PD), on the other hand, originates from a different tradition and focuses on involving stakeholders—especially end users or marginalized groups—directly in the design process. The core idea of PD is to democratize design decision-making and ensure outcomes reflect the needs and values of those affected (Simonsen and Robertson, 2012). While classic PD was applied to developing real systems, there has been growing interest in participatory speculative design: inviting participants to co-create visions of the future. This fusion aims to overcome one criticism of early speculative design and design fiction: that they sometimes reflected the narrow, “privileged” perspectives of their authors (Tonkinwise, 2014). Scholars have pointed out that many speculative design works, especially from Western academic or art contexts, carried assumptions not shared by broader populations, and risked a kind of cultural or technological elitism. In response, methods like participatory design fiction (e.g. workshops where diverse groups build fictional scenarios together) and speculative co-design have been developed to inject plurality and ground speculation in a wider range of experiences (Lindley et al., 2014; Gerber et al., 2018).
A systematic literature review on participatory speculative design identifies this practice as shifting focus from artifacts to process, empowering public participation across multiple stakeholders to “co-create potential pluralistic futures and democratize imagination” (Zhang, 2024). This research outlines three practical pathways—technical speculation, social speculation, and integrated speculation—along with seven participatory methods and a four-phase framework to guide participatory speculative design practice.
2.2.3. From Artifacts to Infrastructures
@@
Another notable evolution in design fiction practice has been a shift from focusing on single artifacts to considering systems and infrastructures. Sterling’s notion of keeping design fiction “on the level of objects and services” was a useful starting heuristic, but many contemporary design researchers have sought to reconnect those isolated artifacts to the larger contexts that give them meaning. As Wong et al. (2020) observe, speculative designs “often imply a lifeworld” even if they only show a gadget. There is an increasing push to explicate that lifeworld: to design not just a speculative object but also illustrate the networks, practices, regulations, and cultures around it. Infrastructural Speculations (Wong et al., 2020) exemplify this trend. They outline design tactics to deliberately incorporate infrastructure considerations into speculative design, so that the speculation “places lifeworlds at the center of design concern” and highlights the cultural, regulatory, environmental, and repair conditions enabling the imagined future. Their work suggests techniques like “infrastructural inversion”—essentially flipping the focus to the normally invisible support systems—and considering multiple stakeholders and temporal scales in speculative scenarios.
Wakkary et al. (2015) propose material speculation as a complementary approach that “utilizes actual and situated design artifacts in the everyday as a site of critical inquiry.” Rather than relying solely on scenarios, prototypes, and forecasting, material speculation emphasizes using real, tangible artifacts to provoke critical reflection by coupling counterfactual artifacts—objects designed to challenge everyday logic—with literary concepts of possible worlds.
@@ 2.3. Infrastructure Studies, Temporality, and Sociotechnical Imaginaries
Our work is additionally informed by the rich body of infrastructure studies in both HCI and Science & Technology Studies (STS). Infrastructure in this context refers to the large-scale sociotechnical systems and underlying structures that support human activities—from transportation grids and power networks to information systems and organizational routines. A foundational perspective by Star and Ruhleder (1996) reconceptualized infrastructure not as a static thing, but as something that is relational, ecological, and learned through use. Bowker et al. (2010) argue for “information infrastructure studies” as a way of attending to the mundane yet consequential work of standardization, classification, and data curation that underpins contemporary knowledge practices. Edwards (2003) outlines an agenda for infrastructure studies that emphasizes heterogeneity, embeddedness, and the long temporal horizons of infrastructural systems. Ribes and Finholt (2007) describe the long-term planning of infrastructure development, highlighting tensions between innovation and stability, local flexibility and global interoperability, short-term projects and long-term maintenance.
A recent systematic literature review by Lyu et al. (2025) identifies three major themes in SIGCHI infrastructure research:
@@
Subsequent work has elaborated concepts such as infrastructure time and infrastructuring. Karasti and Baker (2004) examine how collaborative data infrastructures in science require ongoing care, negotiation, and reconfiguration over years and decades, rather than one-off design interventions. Infrastructuring, drawn from participatory design, foregrounds users and communities as co-designers who continuously shape, appropriate, and re-align infrastructures in practice (Karasti and Baker, 2004; Björgvinsson et al., 2012). Larkin (2013) analyzes “the politics and poetics of infrastructure”, showing how infrastructures materialize state power, cultural imaginaries, and sensory experiences. Anthropological accounts of infrastructure in cities, energy systems, and logistics further emphasize uneven access, differential vulnerabilities, and the improvisational practices people develop to cope with breakdowns (Anand et al., 2018; Harvey, 2012).
Crucially for our purposes, STS has also foregrounded temporality and sociotechnical imaginaries as central to infrastructural politics. Jasanoff and Kim (2015) conceptualize sociotechnical imaginaries as collectively held, institutionally stabilized visions of desirable futures that guide large-scale techno-political projects. Sareen et al. (2021) argue that studying energy infrastructures requires explication of temporality, including how expectations, delays, and irreversibilities shape what futures become possible. Infrastructure studies thus provide both conceptual tools (e.g., inversion, infrastructuring) and temporal sensibilities (e.g., long-now, path-dependence, legacy) that are highly relevant to protocol design.
@@ 3. Methodological Framework: Protocol Futuring
Protocol Futuring is a methodological framework for speculating on the future of sociotechnical infrastructures by using design fiction techniques centered on protocols. In essence, it asks researchers, designers, and participants to imagine future worlds where the key drivers or points of inquiry are protocols—whether technical standards, social rules, or political processes—and to explore the second-order consequences of those protocols in action. This approach extends speculative design into what we call an “infrastructural lifeworld” mode, where the unit of analysis isn’t a single artifact or user experience but rather the network of relationships, agreements, and hidden work that a protocol coordinates or disrupts. This situates our method within, but distinct from, established speculative design traditions (Star and Ruhleder, 1996; Bowker and Star, 1999; Pipek and Wulf, 2009).
@@
In contrast to previous work, our primary aim is to shift speculative design’s focus from products to protocols. We concentrate on new protocols and how they function in future lifeworlds. Rather than emphasizing specific individuals or heroes in our speculative fiction, we focus on the systemic effects of protocols. Our approach is guided by what we call Chiang’s Law, named after science fiction author Ted Chiang (Rao, 2025): “science fiction is about strange rules, while fantasy is about special people.” This aligns the method with recent moves in HCI to examine governance-by-infrastructure, standards, and rule systems (Lampland and Star, 2009; Ribes, 2017), yet uses fiction to expose how those rules behave over extended temporal horizons.
@@
A primary aim is to move beyond first-order effects (“new technology X performs function Y”) to examine second-order and longer-term effects (as X becomes widespread, it causes Z unintended outcome among society/environment, or actors respond with W adaptation). Second-order effects are explored not as deterministic predictions but as plausible dynamics grounded in what Edwards (2003) calls the ”vast machine” of infrastructural interdependence. The method deliberately directs attention to what happens after a protocol’s initial deployment and uptake—essentially, to consider the trajectory of a protocol through time (e.g., years, decades) and scale (from small pilot to global infrastructure). This exploration includes phenomena like protocol drift (implementations or uses diverging from original design over time), protocol jams (congestions or deadlocks in socio-technical systems caused by interacting protocols, similar to traffic jams), and protocol ossification (protocols becoming rigid and resistant to change as they entrench). For example, a Protocol Futuring about a future internet standard might explore how that standard initially improves security but later leads to a new form of digital lock-in or black market to circumvent its restrictions—outcomes beyond the protocol’s immediate goal. By targeting these ripple effects, Protocol Futuring aligns with the ethos of “predicting the traffic jam, not just the car” (Pohl, 1968), using imaginative narrative as a foresight tool. It is worth noting that here, we treat drift, jam, and ossification not as predictions or diagnoses, but as interpretive hypotheses generated through speculative enactment. These dynamics function as candidate patterns that invite later empirical investigation rather than claims about inevitable outcomes.
“A good science fiction story should be able to predict not the automobile but the traffic jam.” – Frederik Pohl
@@
In Protocol Futuring, context is not backdrop but a primary analytic unit: scenarios are constructed to show how a protocol operates within—and is reshaped by—the social norms, institutions, regulatory conditions, and legacy infrastructures that constitute its lifeworld. This aligns the method with infrastructural speculation (Wong et al., 2020) and with feminist and care-centered infrastructuring traditions that foreground ongoing maintenance and situated negotiation (Pipek and Wulf, 2009; Karasti, 2014). Rather than presenting protocols as neutral technical rules, Protocol Futuring depicts how they are interpreted, contested, and maintained over time. This may involve narrative techniques such as following invisible actors (e.g., maintainers, governance committees) or describing scenes from the protocol’s point of view. The aim is to make the infrastructural “wires and threads” visible and to situate protocols within the lived practices that give them force.
@@
Because protocols distribute authority across designers, implementers, users, regulators, adversaries, and bystanders, Protocol Futuring relies on participatory methods to surface these divergent viewpoints. Building on traditions in participatory design (Ehn, 2008; Bjögvinsson et al., 2012) and game-based co-design (Gaver et al., 1999; Brandt et al., 2012), our approach invites multiple contributors to co-construct speculative scenarios through workshops, role-play sessions, and open fiction calls. This distributes authorship beyond designers and enables participants to inhabit different stakeholder positions—everyday users, maintainers, marginalized groups, or adversaries—thereby generating plural, sometimes conflicting, worldviews. Such plurality aligns with the need to imagine protocols that require collective adoption and negotiation rather than heroic individual use. In practice, we draw on PD facilitation techniques (e.g., scenario mapping) and design-game mechanics (Brandt et al., 2012) to structure interaction. These participatory and game-like structures make the infrastructural politics of protocols experientially accessible and foreground how rules shape—and are shaped by—collective imagination.
@@
Although protocols can seem abstract, Protocol Futuring materializes speculative worlds through concrete artifacts, drawing from traditions in critical design (Dunne and Raby, 2013), design fiction (Sterling, 2005), and future-oriented prototyping (Halse et al., 2010). Outputs may include short stories, fictional policy documents, speculative diagrams, interface mockups, or immersive videos—forms that make the protocol’s logic, breakdowns, and sociopolitical effects legible. These artifacts function not as “hero objects” but as probes that reveal how a rule system organizes everyday life. For instance, a fabricated citizen handbook for a future data protocol or a visualization of rerouted network flows during a protocol split can foreground operational constraints, power dynamics, and maintenance work. We often employ cross-media storytelling—text, images, interactive maps, AI-generated video—to surface different infrastructural layers. As demonstrated in the South Beast Asia case presented in the Appendix, such multimodal artifacts help participants inhabit the protocol’s world and expose second-order dynamics that would remain invisible in purely textual speculation. Protocol Futuring thus describes not a fixed output format but a strategic orientation capable of using any medium necessary to render protocol-driven futures thinkable.
@@
Methodologically, we treat Protocol Futuring as a form of research-through-design (RtD) (Blythe, 2014). This means the act of creating the fiction is itself a mode of inquiry: by designing these speculative scenarios, we are effectively probing research questions about protocols (“what if this protocol were in place, how would X or Y play out?”) and generating knowledge in the form of narrative insights, identified pitfalls, or design considerations, following Zimmerman et al. (2007). Throughout our process, we maintain a reflexive stance—we document the decision-making, the alternatives considered, the influences from real-world references, and the participants’ reactions. Analysis of a protocol fiction may involve looking for emergent themes across different stories, or using the fiction as a provocation in interviews with experts. For example, after producing a set of stories in the Knowledge Futurama workshop, we conducted a debrief discussion where librarians and archivists (some of whom participated) reflected on the scenarios: which seemed plausible, which touched on real concerns, which surprised them by revealing a blind spot in current practices. This reflective step, aligning with Ghajargar and Bardzell (2021), ensures that the fictional outputs are brought back to bear on present understanding, closing the loop of speculation to inform reality.